OpenStudy (aravindg):

Let alpha and beta be the solutions of the equation f(x)=0,where f(x)=p*cos x+q*sin x,show that f(x) vanishes identically if alpha-beta not equal to n*pi,n belongs to integers.

OpenStudy (aravindg):

help pls

OpenStudy (aravindg):

yey

OpenStudy (jamesj):

First we need a trig identity. Consider the dot product of two vector (p,q) and (cos x, sin x). Note that (p,q) makes an angle to positive x-axis, phi = arctan(q/p). Then\[(p,q).(\cos x, \sin x) = p \cos x + q \sin x\] also \[(p,q).(\cos x, \sin x) = \sqrt{p^2 + q^2} \cos (x - \phi)\] So the two expressions are equal.

OpenStudy (jamesj):

Suppose now that alpha and beta are zeros of f(x)=p*cos x+q*sin x i.e., zeros of \[f(x) = \sqrt{p^2 + q^2} \cos(x - \phi)\] The cos part of this function, cos(x - phi) has zeros at x = k.pi - phi for integers k

OpenStudy (jamesj):

thus if \[\alpha - \beta \neq k \pi \text{ , for some integer k}\] Then at least one of alpha and beta are not zeros of cos(x - phi) Hence \[\sqrt{p^2 + q^2} = 0 \ \ \ \text{ i.e., } p = q = 0\]

OpenStudy (jamesj):

CORRECTION The function cos(x - phi) has zeros at \[x = k \pi + \phi \text{, for integers } k\]